# Efficiently and Effectively: A Two-stage Approach to Balance Plaintext and Encrypted Text for Traffic Classification

### v2

@@ Abstract.
Encrypted traffic classification is the task of identifying the application or service associated with encrypted network traffic. One effective approach for this task is to use deep learning methods to encode the raw traffic bytes directly and automatically extract features for classification (byte-based models). However, current byte-based models input raw traffic bytes, whether plaintext or encrypted text, for automated feature extraction, neglecting the distinct impacts of plaintext and encrypted text on downstream tasks. Additionally, these models primarily focus on improving classification accuracy, with little emphasis on the efficiency of models. In this paper, for the first time, we analyze the impact of plaintext and encrypted text on the model’s effectiveness and efficiency. Based on our observations and findings, we propose a two-phase approach to balance the trade-off between plaintext and encrypted text in traffic classification. Specifically, Stage one is to Determine whether the Plain text is enough to be accurately Classified (DPC) using the proposed DPC Selector. This stage quickly identifies samples that can be classified using plaintext, leveraging explicit byte features in plaintext to enhance model’s efficiency. Stage two aims to adaptively make a classification with the result from stage one. This stage incorporates encrypted text information for samples that cannot be classified using plaintext alone, ensuring the model’s effectiveness on traffic classification tasks. Experiments on two datasets demonstrate that our proposed model achieves state-of-the-art results in both effectiveness and efficiency.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Encrypted traffic classification is a critical task in the field of cybersecurity, focusing on identifying and categorizing network traffic data that has been encrypted (Bujlow et al., 2015; Aceto et al., 2020). This process is essential for maintaining network security, managing bandwidth, detecting malicious activities, and ensuring regulatory compliance (Akbari et al., 2021; Cui et al., 2023). Applications of encrypted traffic classification span various domains, including intrusion detection systems, quality of service (QoS) management, and network forensics. As illustrated in Figure 1, some common applications include encrypted application classification (Lin et al., 2022), attack detection (Meng et al., 2021), VPN classification (Draper-Gil et al., 2016), etc.
@@
Figure 1. An example of traffic data packet. The right-hand part means the application classification task. Green font represents the different categories.
Figure 2. Testing time and accuracy analysis on different settings between plaintest and encrypted text on four datasets.
@@
Network traffic is composed of packets, as shown in Figure 1, each consisting of two main components: the header (plaintext) and the payload (usually encrypted text). The header contains metadata about the packet, such as source and destination IP addresses, protocol information, and so on. The payload, on the other hand, contains the actual data being transmitted and is often encrypted in the form of hexadecimal to ensure confidentiality and integrity. Current byte-based encrypted traffic classification researches typically use both the header and the payload as inputs to the models. A fixed number of bytes, like 128, 256, etc., are utilized for the classification task. These bytes often include a mix of plaintext information from the header and encrypted data from the payload. At the same time, studies (Akbari et al., 2022; Zhang et al., 2023) have shown that plaintext information significantly contributes to the effectiveness of classification. For example, the paper (Akbari et al., 2022) mentions that the availability of some obvious feature of the server (e.g., plain-text SNI field) in-the-clear is crucial for the utility of deep-learning-based (DL) approaches. Another research (Zhang et al., 2023) proves through ablation experiments that plaintext information is more helpful for the accuracy of classification effect than encrypted text information. This leads to a fascinating hypothesis: Can we achieve satisfactory traffic classification performance using only the plaintext information present in the header?
@@
To further validate the roles of plaintext, encrypted text, and their combination in traffic classification, we compare the one of the state-of-the-art (SOTA) models (Lin et al., 2022; Zhao et al., 2023) performance and time overhead across four datasets. For the combined plaintext-encrypted text setting, as shown in Figure 2(b), the current prevalent methods that utilize both plaintext and encrypted text achieve the highest accuracy (indicated by the blue line). However, this setting results in high time overhead, as depicted in Figure 2(a).
@@
Some studies have demonstrate that the effectiveness of DL-based methods largely rely on plaintext information. This suggests that satisfactory classification performance can be achieved by leveraging only plaintext data. Notably, for the plaintext-only setting, as indicated by the gray line in Figure 2(b), the model achieves around 90% accuracy on certain datasets, further substantiating the crucial role of plaintext in classification. Surprisingly, this setting also enhances time efficiency by approximately 4 to 10 times. However, relying solely on plaintext data obtains lower accuracy on some situations.
@@
In conclusion, the time overhead for the combined plaintext-encrypted text setting is considerably high, although the accuracy is the best in these settings. Therefore, balancing the utilization of plaintext and encrypted text to ensure accurate classification while optimizing time efficiency remains a key research challenge. Our study aims to explore this trade-off to achieve both high accuracy and minimal time overhead in encrypted traffic classification.
To address the above issues, we propose a simple but Efficient and Effective Two-Phase approach (EETP) to balance plaintext and encrypted text for traffic classification. In the first stage, we introduce a new traffic classification task: Determining whether the Plaintext is enough to be accurately Classified (DPC) and propose a DPC selector. The purpose of this stage is to fully leverage plaintext information to quickly and efficiently determine if the traffic can be classified using only plaintext. This allows us to optimize the utilization of time by avoiding unnecessary processing of encrypted data when plaintext alone suffices. In the second stage, we proceed with the plaintext classification model for a efficient classification if the traffic can be classified only using plaintext. Otherwise, a combination of plaintext and encrypted text is considered to enhance the model’s accuracy with the encrypted text classification model. The two-phase design ensures a balance between prediction accuracy and efficiency. By attempting to classify with plaintext, DL- based models can significantly reduce processing time for a large portion of the traffic. For cases where plaintext is insufficient, incorporating encrypted text in the second phase ensures that models still achieve high accuracy as well as low time consumption.
@@
  • For the first time, we analyze the impact of plaintext and encrypted text on model performance and time overhead. We introduce an interesting idea of categorizing data packets into two types: plaintext classifiable and plaintext non-classifiable. Based on this categorization, a novel approach is designed to significantly reduce computational overhead while maintaining high prediction accuracy.
@@
  • We propose a two-phase approach to balance the trade-off between plaintext and encrypted text in traffic classification. Stage one is to Determine whether the Plain text is enough to be accurately Classfied (DPC) with the proposed DPC Selector. Stage two aims to adaptively make a classification with the result from stage one.
@@
  • Experimental results on one public dataset and one real-world dataset collected by ourselves demonstrate the efficiency and effectiveness of our approach.
@@
Figure 3. The overview of EETP, which consists of two stages. Stage one is to determine whether the Plain text is enough to be accurately classfied with the proposed DPC Selector. Stage two is to adaptively make a classification with the result from stage one. The green arrows indicate the plaintext processing, brown arrows present the encrypted text processing.
@@ 2. Models
The proposed approach EETP is presented in Figure 3, which efficiently and effectively to balance plaintext and encrypted text for traffic classification. The core modules consist of a DPC selector (in stage one) and plaintext or encrypted text classification models (in stage two). Specifically, the DPC selector is able to determining whether the plaintext is enough to be accurately classified. The purpose of this module is to fully leverage plaintext information to quickly and efficiently determine if the traffic can be classified using only plaintext. Then, plaintext or encrypted text classification models can make an adaptively classification with the result from stage one. In the following, we first describe the problem formulation. Then, the detail of the EETP is introduced step by step.
@@ 2.1. Problem Formulation
Given a pcap file, the scapy toolkit parses the file and extracts the network packets. Each packet includes header portion (plaintext) $X^{p}$ and payload data (encrypted text) $X^{e}$ in the form of string $X=\{X^{p},X^{e}\}=\{x^{p}_{1},x^{p}_{2},\dots,x^{p}_{M},\dots,x^{e}_{N}\}$ that consists of ${N}$ traffic words, with a category label $y$, e.g. YouTube, Facebook, etc., where $p$ indicates plaintext, $e$ indicates encrypted text, $M$ means the number of plaintext traffic words. Our task can be divided into two subtasks. The first subtask involves determining, based on plaintext information, whether the data is sufficient for accurate classification. This is a binary classification task. The second subtask depends on the result of the first subtask, selecting different branches to output the corresponding packet category label, which is a multi-class classification task.
@@ 2.2.1. DPC Data Construction
In the first stage of training, the proposed new DPC task involves assessing whether the plaintext information is sufficient for accurate classification. However, existing datasets do not contain labels for this specific purpose, making the construction of the DPC data a crucial step.
Considering that the DPC task is a binary classification problem and relatively straightforward, many SOTA traffic classification models (MengWMLLZ22; Lin et al., 2022) can be leveraged for data labeling. Specifically, we use only plaintext information as input and employ the Pacrep model to perform standard traffic classification training. Once the model is trained, we predict the labels for all data samples. A data sample is labeled as “1” (plaintext classifiable) only if the predicted label matches the original sample’s label. Otherwise, the sample is labeled as “0” (plaintext non-classifiable). The formulation can be defined as follows:
@@
After obtaining all the labels, we can then train the DPC Selector, which will enable us to distinguish between classifiable and non-classifiable plaintext data effectively. This process ensures accurate and efficient traffic classification in subsequent stages.
Table 1. Comparison results on public and collected datasets. indicates reproduced results. Time metric means that the number of samples that the model predicts per second.
Methods ISCXVPN CHNAPP
Macro-F1 Micro-F1 Time/ each second Macro-F1 Micro-F1 Time/ each second
APPS (Taylor et al., 2018) 81.68 81.33 - - - -
SMT (Shen et al., 2019) 76.31 76.00 - - - -
Deep Packet (Lotfollahi et al., 2020) 65.06 66.70 - - - -
TR-IDS (Min et al., 2018) 54.98 57.66 - - - -
HEDGE (Casino et al., 2019) 29.67 34.00 - - - -
3D-CNN (Zhang et al., 2020) 54.34 57.00 - - - -
BERT (Devlin et al., 2019) 88.45 88.73 - 56.94 84.08 -
PacRep (Meng et al., 2022) 99.32 99.13 22.75 95.10 98.75 44.46
ET-BERT (Lin et al., 2022) 94.25 96.05 - 94.08 96.09 -
YaTC (Zhao et al., 2023) 96.67 98.04 - 90.44 93.50 -
EEPT (Ours) 99.89 99.97 124.01 94.95 99.14 82.11
@@ 2.2.2. DPC Selector
After obtaining the training data, we proceed to train the DPC selector. This module aims to filter traffic data, directing plaintext classifiable samples to the plaintext classification model and plaintext non-classifiable samples to the encrypted classification model. This adaptive approach enhances the effectiveness of classification for different types of data, as well as playing a crucial role in optimizing the efficiency of traffic classification. Specifically, existing off-the-shelf traffic classification model  [^1] https://github.com/ict-net/PacRep is utilized as the DPC selector to ensure robust performance. The cross-entropy loss function is used for training. Then, we test the model to selectively separate the samples in the test set into two categories. As illustrated in Figure 3, our experiments show that the majority of data (represented by the green portion) is directed towards the plaintext classification model, while a smaller portion is directed towards the encrypted text classification model. This demonstrates the superiority of our proposed framework, effectively and efficiently achieving traffic classification.

### v3

@@ Abstract.
Encrypted traffic classification is the task of identifying the application or service associated with encrypted network traffic. One effective approach for this task is to use deep learning methods to encode the raw traffic bytes directly and automatically extract features for classification (byte-based models). However, current byte-based models input raw traffic bytes, whether plaintext or encrypted text, for automated feature extraction, neglecting the distinct impacts of plaintext and encrypted text on downstream tasks. Additionally, these models primarily focus on improving classification accuracy, with little emphasis on the efficiency of models. In this paper, for the first time, we analyze the impact of plaintext and encrypted text on the model’s effectiveness and efficiency. Based on our observations and findings, we propose an efficient and effective two-stage approach to balance the trade-off between plaintext and encrypted text in traffic classification. Specifically, stage one proposes a DPC selector to Determine whether the Plaintext information is sufficient to perform subsequent Classification (DPC). This stage quickly identifies samples that can be classified using plaintext, leveraging explicit byte features in plaintext to enhance model’s efficiency. Stage two aims to adaptively make a classification with the result from stage one. This stage could incorporate encrypted text information for samples that cannot be classified using plaintext alone, ensuring the model’s effectiveness on traffic classification tasks. Experiments on two public datasets and one real-world collected dataset demonstrate that our proposed model achieves state-of-the-art results in both effectiveness and efficiency. The model code will be released in https://github/after/review.
Traffic classification, Plain and encrypted text, Model efficiency
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Security and privacy Network securityccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Encrypted traffic classification is a critical task in the field of cybersecurity, focusing on identifying and categorizing network traffic data that has been encrypted (Bujlow et al., 2015; Aceto et al., 2020). This process is essential for maintaining network security, managing bandwidth, detecting malicious activities, and ensuring regulatory compliance (Akbari et al., 2021; Cui et al., 2023). Applications of encrypted traffic classification span various domains, including intrusion detection systems, quality of service (QoS) management, and network forensics. As illustrated in Figure 1, some common applications include encrypted application classification (Lin et al., 2022), attack detection (Meng et al., 2021), VPN classification (Draper-Gil et al., 2016; Oh et al., 2023), etc.
@@
Figure 1. An example of traffic data packet. The right-hand part means the downstream traffic classification task. Green font represents the different labels.
@@
Network traffic is composed of packets, as shown in Figure 1, each consisting of two main components: the header (plaintext) and the payload (usually encrypted text). The header contains metadata about the packet, such as source and destination IP addresses, protocol information, and so on. The payload, on the other hand, contains the actual data being transmitted and is often encrypted in the form of hexadecimal to ensure confidentiality and integrity. Current byte-based encrypted traffic classification researches (Liu et al., 2019; Jemal et al., 2021; Meng et al., 2022; Zhao et al., 2023) typically use both the header and the payload as inputs to the models. A fixed number of bytes, like 128, 256, 512, etc., is utilized for the classification task. These bytes often include a mix of plaintext information from the header and encrypted data from the payload. At the same time, some studies (Akbari et al., 2022; Zhang et al., 2023) have shown that plaintext information significantly contributes to the effectiveness of classification. For example, the paper (Akbari et al., 2022) mentions that the availability of some obvious feature of the server (e.g., plain-text SNI field) in-the-clear is crucial for the utility of deep-learning-based (DL) approaches. Another research (Zhang et al., 2023) proves through ablation experiments that plaintext information is more helpful for the effectiveness of classification than encrypted text information. This leads to a fascinating hypothesis: Can we achieve satisfactory traffic classification performance using only the plaintext information present in the header?
Figure 2. Testing time and accuracy analysis on different settings between plaintext and encrypted text on four datasets.
@@
To further validate the roles of plaintext, encrypted text, and their combination in traffic classification, we compare the one of the state-of-the-art (SOTA) models (Meng et al., 2022) performance and time overhead across four datasets (Draper-Gil et al., 2016; Lashkari et al., 2017; Wang et al., 2017). For the combined plaintext-encrypted text setting, as shown in Figure 2(b), the current prevalent methods that utilize both plaintext and encrypted text achieve the highest accuracy (indicated by the blue line). However, this setting results in high time overhead, as depicted in Figure 2(a) blue line.
@@
Some studies (Akbari et al., 2022; Zhang et al., 2023) have demonstrate that the effectiveness of DL-based methods largely rely on plaintext information. This suggests that satisfactory classification performance can be achieved by leveraging only plaintext data. Notably, for the plaintext-only setting, as indicated by the gray line in Figure 2(b), the model achieves around 90% accuracy on certain datasets, further substantiating the crucial role of plaintext in classification. Surprisingly, this setting also enhances time efficiency by approximately 4 to 10 times (indicated by the gray line in Figure 2(a)). However, relying solely on plaintext data obtains lower accuracy on some situations, like result on USTC and CHNAPP datasets (Wang et al., 2017) in Figure 2(b).
@@
In conclusion, the time overhead for the combined plaintext-encrypted text setting is considerably high, although the accuracy is the best in these settings. Therefore, balancing the utilization of plaintext and encrypted text to ensure accurate classification while optimizing time efficiency remains a key research challenge. Our study aims to explore this trade-off to achieve both high accuracy and low time overhead in traffic classification.
To address the above issues, we propose a simple but Efficient and Effective Two-Stage approach (EETS) to balance plaintext and encrypted text for traffic classification. In the first stage, we introduce a new traffic classification task: Determining whether the Plaintext is sufficient to perform subsequent Classification (DPC) and propose a DPC selector. The purpose of this stage is to fully leverage plaintext information to quickly determine if the traffic can be classified using only plaintext (plaintext classifiable). This allows us to enhance the time efficiency by avoiding unnecessary processing of encrypted data when plaintext alone suffices. In the second stage, we proceed with the plaintext classification model for a efficient classification if the traffic can be classified only using plaintext. Otherwise (plaintext non-classifiable), a combination of plaintext and encrypted text is considered to further improve the model’s accuracy. The two-stage design ensures a balance between effectiveness and efficiency. By attempting to classify with plaintext, DL-based models can significantly reduce processing time for a large portion of the traffic. For cases where plaintext is insufficient, incorporating encrypted text ensures that models still achieve high accuracy, while maintaining high time efficiency.
@@
  • For the first time, we analyze the impact of plaintext and encrypted text on model performance and time overhead. Based on this, we introduce an interesting idea of categorizing data packets into two types: plaintext classifiable and plaintext non-classifiable. Then, a simple but efficient and effective approach is designed to significantly reduce time overhead while maintaining high classification accuracy.
@@
  • We propose a two-stage approach to balance the trade-off between plaintext and encrypted text in traffic classification. Stage one proposes a DPC selector to determine whether the plaintext information is sufficient to perform subsequent classification. Stage two is to adaptively make a classification with the result from stage one.
@@
  • Experimental results on two public datasets and one real-world dataset collected by ourselves demonstrate the efficiency and effectiveness of our approach.
@@
Figure 3. The overview of EETS, which consists of two stages. Stage one proposes a DPC selector to determine whether the plaintext is sufficient to perform subsequent classification. Stage two is to adaptively make a classification with the result from stage one. The green arrows indicate the plaintext processing, red arrows present the plaintext and encrypted text processing.
2. Related Work
2.1. Plaintext Classification
In the early days of network traffic classification, methods (Cheng and Wang, 2011; Yoon et al., 2009; Zhang et al., 2014; Shbair et al., 2015, 2016) mainly rely on plaintext information to analyze traffic data. This includes using transport layer protocol port numbers, such as UDP or TCP, or leveraging plaintext information like the Server Name Indication (SNI). Port-based methods (Cheng and Wang, 2011; Yoon et al., 2009; Zhang et al., 2014) are straightforward to implement and have low time complexity, making them a popular choice when only specific port applications needed to be classified. However, as applications and protocols diversified and techniques such as port hopping and port masquerading emerge, the accuracy of port-based classification methods decrease, rendering them unreliable. SNI-based methods (Shbair et al., 2015, 2016) use the hostname information requested during the TLS handshake to determine the traffic category. These methods boast high accuracy and are easy to implement. Nevertheless, not all clients support SNI, particularly older browsers or operating systems. When such clients attempt to access SNI-enabled websites, SNI-based methods fail.
Rencently, with the continuous evolution of encryption protocols (Han et al., 2020; Thakur et al., 2023), plaintext classification models have become increasingly limited. Early plaintext classification approaches heavily relied on manually crafted feature engineering, and solely using plaintext information may not suffice for certain scenarios. As a result, researchers have shifted towards DL-based models, which allow for automatic feature extraction without the need to restrict the data to plaintext or encrypted text. This approach enhances the adaptability and robustness of network traffic classification methods.
2.2. Encrypted Traffic Classification
From a methodological perspective, encrypted traffic classification techniques can be broadly categorized into feature-based methods and byte-based models. Given that this paper directly uses both the header and payload content as inputs, the focus will be on byte-based models (Sirinam et al., 2018; Liu et al., 2019; Jemal et al., 2021; Meng et al., 2022; Zhao et al., 2023). The details are as follows.
Byte-based models leverage raw bytes sequences of network traffic for classification. These models do not rely on manually crafted features but utilize the inherent information present in the raw bytes. This approach allows the models to automatically learn the relevant patterns and features necessary for accurate classification, making them highly suitable for handling encrypted traffic where traditional feature extraction methods fall short.
Byte-based models (Sirinam et al., 2018; Liu et al., 2019; Jemal et al., 2021; Meng et al., 2022; Zhao et al., 2023) work directly on raw traffic bytes and learn representations using techniques like Convolutional Neural Networks (CNNs) and Long Short-Term Memory Networks (LSTMs). They can capture complex patterns in encrypted data packets without manual feature engineering. These approaches can be categorized as image-based or text-based depending on how they interpret the input traffic. Specifically, image-based models (Sirinam et al., 2018; Jemal et al., 2021; Zhao et al., 2023) treat the traffic data as 2D images by transforming the raw bytes into grayscale intensity values or RGB colors. They then apply well-known image classification CNN architectures like ResNet and Inception that have shown excellent performance in computer vision. These models can identify visual patterns in the transformed input matrix based on learned convolutional filters. Text-based models such as Recurrent Neural Networks (RNN) and LSTMs process the input sequentially as text sequences (Lotfollahi et al., 2020; Meng et al., 2022). They can analyze long-range dependencies in the raw traffic bytes. Self-attention models like transformer are also popular for learning contextual relationships in encrypted traffic data. For example, Meng et al. (2022) utilize contrastive loss with a sample selector and attention mechanism to optimize the learned representations with various traffic classification tasks.
@@ 3. Models
The proposed approach EETS is presented in Figure 3, which efficiently and effectively to balance plaintext and encrypted text for traffic classification. The core modules consist of a DPC selector (in stage one) and plaintext or encrypted text classification models (in stage two). Specifically, the DPC selector is able to determine whether the plaintext information is sufficient to perform subsequent classification. The purpose of this module is to fully leverage plaintext information to quickly determine if the traffic can be classified using only plaintext. Then, plaintext or encrypted text classification models can make an adaptively classification with the result from stage one. In the following, we first describe the problem formulation. Then, the detail of the EETS is introduced.
@@ 3.1. Problem Formulation
Given a pcap file, the scapy toolkit parses the file and extracts the network packets. Each packet includes header portion (plaintext) $X^{p}$ and payload data (usually encrypted text) $X^{e}$ in the form of string $X=\{X^{p},X^{e}\}=\{x^{p}_{1},x^{p}_{2},\dots,x^{p}_{M},\dots,x^{e}_{N}\}$ that consists of ${N}$ traffic words, with a category label $y$, e.g. YouTube, Facebook, etc., where $p$ indicates plaintext, $e$ indicates encrypted text, $M$ means the number of plaintext traffic words. Our task can be divided into two subtasks. The first subtask involves determining, based on plaintext information, whether the data is sufficient for subsequent classification. The second subtask depends on the result of the first subtask, selecting different branches to output the corresponding packet category label, which is a multi-class classification task.
@@ 3.2.1. DPC Data Construction
In the first stage of training, the proposed new DPC task involves assessing whether the plaintext information is sufficient for subsequent classification. However, existing datasets do not contain labels for this specific purpose, making the construction of the DPC data a crucial step.
Considering that the DPC task is a traffic classification problem and relatively straightforward, many SOTA traffic classification models (Lin et al., 2022; Meng et al., 2022) can be leveraged for data labeling. Specifically, we use only plaintext information as input and employ the Pacrep model to perform standard traffic classification training. Once the model is trained, we predict the labels for all data samples. A data sample is labeled as “1” (plaintext classifiable) only if the predicted label matches the original sample’s label. Otherwise, the sample is labeled as “0” (plaintext non-classifiable). The formulation can be defined as follows:
@@
After obtaining all the labels, we can then train the DPC selector, which will enable us to distinguish between plaintext classifiable and non-classifiable effectively. This process ensures accurate and efficient traffic classification in subsequent stage.
@@ 3.2.2. DPC Selector
After obtaining the training data, we proceed to train the DPC selector. This module aims to filter traffic data, directing plaintext classifiable samples to the plaintext classification model and plaintext non-classifiable samples to the encrypted classification model. This adaptive approach enhances the effectiveness of classification for different types of data, as well as playing a crucial role in optimizing the efficiency of traffic classification. Specifically, existing off-the-shelf traffic classification model  [^1] https://github.com/ict-net/PacRep is utilized as the DPC selector to ensure robust and effective performance. The cross-entropy loss function is used for training. Then, we test the model to selectively separate the samples in the test set into two categories. As illustrated in Figure 3, our experiments show that the majority of data (represented by the green portion) is directed towards the plaintext classification model, while a smaller portion is directed towards the encrypted text classification model. This demonstrates the superiority of our proposed framework, effectively and efficiently achieving traffic classification.
Figure 4. An overview of plaintext classification model.